# MFC 5.0: An exascale many-physics flow solver

### v1

@@ 4.3.6 Metaprogramming
Metaprogramming is implemented using the Fortran preprocessor Fypp [79]. Fypp is a Python-based preprocessor that is used to inline subroutines via programmer directives and collapse repetitive code. MFC uses Fypp for several tasks, including automation of deep copies of derived types for allocation on the GPU device. NVIDIA’s NVHPC Fortran compiler automatically allocates derived types on the device, but CCE adheres closer to the OpenACC technical specification (3.2 at the time of writing) and requires manual allocation. If one holds many host variables, this process is cumbersome. Instead, we use the code in LABEL:lst:deepCopy to inline the source code via @:ACC_SETUP_VFs(var1,var2,…,varN) after allocating the vector fields var1,var2,…,varN on the host. This macro performs deep copies, reducing the number of lines required to initialize the variables at runtime.
@@
Metaprogramming sets problem parameters as constant parameters in case files. This option is called case optimization, enabled via the flag --case-optimization at run time. Case optimization reduces register pressure by providing the compiler with the parameters of a configuration. On CPUs, specifying the problem parameters at compile-time results in a two-times speedup of the code. On GPUs, case optimization decreases run time by about a factor of ten. Variables are precriped in a file with the Fypp code #:set var = val. This file is included in the Fortran source code, compile-time-known parameters are declared using the Fypp logic of LABEL:lst:fyppDec.
@@ 5.3 Taylor–Green vortex
Figure 10: The isosurface with zero Q-criterion of a $\mbox{\text{Re}}=1600$, $\mbox{\text{Ma}}=0.1$ Taylor–Green vortex at dimensionless times (a) $t/t_{c}=4$, (b) $t/t_{c}=12$ and (c) $t/t_{c}=20$. The isosurface is colored by the vorticity magnitude, with darker colors corresponding to higher vorticity.
@@ 6 Solvers with some shared capabilities
Other open-source codes for CFD exist. These and their relative trade-offs compared to MFC 5.0 are briefly discussed here. URANOS-2.0 (De Vanna and Baldan [84]) is a modern Fortran code that uses OpenACC for GPU offloading to NVIDIA and AMD GPUs. URANOS-2.0 contains some of MFC’s shock-capturing capabilities; it also has thorough models for turbulence and focuses on this capability in the context of compressible flow. STREAmS-2 (Sathyanarayana et al. [85]) is also a Fortran code for compressible flow and uses CUDA Fortran and hipFORT to offload to NVIDIA and AMD GPUs. There is also an OpenMP port of STREAmS-2 that makes it portable to Intel GPUs. AFiD-GPU of Zhu et al. [86] is similar in this context. Neko (Jansson et al. [87]) is a particularly modern object-oriented solver that uses multiple levels of abstraction for offloading to GPUs as well as more extensive extensions for vector processors and FPGAs. The above solvers’ offloading capabilities are similar to those of MFC, though the physical models and their sophistication differ strongly from those of MFC. The above solvers are largely used in the context of multi-phase and species flows, for which MFC offers a broad range of modeling capabilities and the numerics that enable their use.
@@ 7 Conclusion
Numerical advancements accompany MFC’s extended feature set. TENO-6 and WENO-Z shock-capturing schemes meaningfully reduce dispersion error compared to conventional WENO schemes. The Strang-split particle solver maintains second-order temporal accuracy for stiff ODE systems. The implementation of a low-Mach preconditioner reduces numerical dissipation for Mach numbers below $0.1$. A ghost-cell immersed boundary method is now available. This method supports complex 2D and 3D geometries that can be imported as level sets or STL files.

### v3

@@ 4.3.6 Metaprogramming
Metaprogramming is implemented using the Fortran preprocessor Fypp [84]. Fypp is a Python-based preprocessor that is used to inline subroutines via programmer directives and collapse repetitive code. MFC uses Fypp for several tasks, including automating deep copies of derived types for allocation on the GPU device. NVIDIA’s NVHPC Fortran compiler automatically allocates derived types on the device, but CCE adheres closer to the OpenACC technical specification (3.2 at the time of writing) and requires manual allocation. If one holds many host variables, this process is cumbersome. Instead, we use the code in LABEL:lst:deepCopy to inline the source code via @:ACC_SETUP_VFs(var1,var2,…,varN) after allocating the vector fields var1,var2,…,varN on the host. This macro performs deep copies, reducing the number of lines required to initialize the variables at runtime.
@@
Metaprogramming sets problem parameters as constant parameters in case files. This option is called case optimization, enabled via the flag --case-optimization at run time. Case optimization reduces register pressure by providing the compiler with the parameters of a configuration. On CPUs, specifying the problem parameters at compile-time results in a two-times speedup of the code. On GPUs, case optimization decreases run time by about a factor of ten. Variables are precriped in a file with the Fypp code #:set var = val. This file is included in the Fortran source code; compile-time-known parameters are declared using the Fypp logic of LABEL:lst:fyppDec.
@@ 5.3 Taylor–Green vortex
Figure 10: The isosurface with zero Q-criterion of a $\mbox{\text{Re}}=1600$, $\mbox{\text{Ma}}=0.1$ Taylor–Green vortex at dimensionless times $t/t_{c}$ as labeled. The isosurface is colored by the vorticity magnitude, with darker colors corresponding to higher vorticity.
5.4 Supersonic gas injection and plume dynamics
Three jets inject into the domain via Dirichlet boundary conditions. The domain has size $[0,16D]\times[0,10D]\times[0,10D]$ where $D$ is the jet diameter and the first dimension is the mean flow direction (horizontal in fig. 11). The domain is discretized by $600\times 400\times 400$ grid cells. The jets have incoming velocity ($V_{\mathrm{jet}}$), pressure, and density that satisfy the Mach 1.5 normal shock conditions; they are organized in an equilateral triangle. The jet Reynolds number is $\mbox{\text{Re}}=\rho V_{\mathrm{jet}}D/\mu=2.5\times 10^{6}$.
Figure 11: A rendering showing the injection and fluid dynamics of three $\mbox{\text{Ma}}=1.5$ jets into quiescent air.
The simulation is executed until dimensionless time $40D/(\mbox{\text{Ma}}\,c)$, where Ma is the Mach number and $c$ is the ambient speed of sound. The last time step is visualized in fig. 11. The visualization shows a volume rendering of the injected fluid after it has meaningfully intruded the domain.
@@ 6 Solvers with some shared capabilities
Other open-source codes for CFD exist. These and their relative trade-offs compared to MFC 5.0 are briefly discussed here. URANOS-2.0 (De Vanna and Baldan [89]) is a modern Fortran code that uses OpenACC for GPU offloading to NVIDIA and AMD GPUs. URANOS-2.0 contains some of MFC’s shock-capturing capabilities; it also has compressible turbulence models. STREAmS-2 (Sathyanarayana et al. [90]) is also a Fortran code for compressible flow and uses CUDA Fortran and hipFORT to offload to NVIDIA and AMD GPUs. There is also an OpenMP port of STREAmS-2 that makes it portable to Intel GPUs. AFiD-GPU of Zhu et al. [91] is similar in this context. Neko (Jansson et al. [92]) is a particularly modern object-oriented solver that uses multiple levels of abstraction for offloading to GPUs as well as more extensive extensions for vector processors and FPGAs. The above solvers have offloading capabilities similar to those of MFC, though the physical models and their sophistication differ. MFC offers a broad range of modeling capabilities and numerics for the broad use of multi-phase and -species flows,
@@ 7 Conclusion
Numerical advancements accompany MFC’s extended feature set. TENO and WENO-Z shock-capturing schemes meaningfully reduce dispersion error compared to conventional WENO schemes. The Strang-split particle solver maintains second-order temporal accuracy for stiff ODE systems. The implementation of a low-Mach preconditioner reduces numerical dissipation for Mach numbers below $0.1$. A ghost-cell immersed boundary method is now available. This method supports complex 2D and 3D geometries that can be imported as level sets or STL files.
@@ Acknowledgements
MRJ acknowledges support from the U.S. Department of Defense under the DEPSCoR program Award No. FA9550-23-1-0485 (PM Dr. Timothy Bentley) and the U.S. National Science Foundation (NSF) under Grant No. 2232427.
AG acknowledges support from the U.S. National Science Foundation (NSF) under Grants CBET 2301721 and CBET 2301709.